# Large Language Model NLP Natural Language Processing

### v1

@@ 3.2 Counterfactual Shifts
Formalisation. We consider common sense contradictions by altering original statements in $c$ into counterfactual ones. The inclusion of counterfactual statements is motivated by the requirement for faithful reasoning. Namely, the validity of a deduction depends only on the structure of the logical form, which in turn should follow the original task description in the context and the question. To introduce counterfactual premises, we do not add new sentences; instead, we negate sentences from the original context $c$. Since original premises often state common sense knowledge, their negation naturally contradicts world knowledge. For example, in the modus ponens inference in Figure 1, we negate $mortal(X)$, stating that All men are immortal. The altered context $c^{\prime}$ is formalized as follows: $c^{\prime}=neg(c)$, where $neg(c)$ negates one of the terms in the original context. The label of the resulting function $f^{\prime}$ is opposite from that of the original function $f$.
Design of counterfactual perturbations. We assume a deductive reasoning dataset where the natural language form corresponds to a logical formula. Then, we introduce counterfactual perturbations by altering each logical formula using predefined rules that negate the formula (see full list of rules in the Appendix). For example, we negate the consequent $\mathsf{q}(a)$ for the first implication of a constructive dilemma and adapt the inference to $\neg\mathsf{q}(a)\vee\mathsf{s}(a)$. Then, the negation to create the context $c^{\prime}$ is manually added before the relevant terms in $c$, thus guaranteeing high data quality. Since the inference results in the opposite label for $f^{\prime}$, we adapt the target label accordingly. Figure 2 (top-right) shows an example of such a contradiction. Importantly, all counterfactual perturbations can be combined with noise perturbations, leading to counterfactual versions of the original ($O_{C}$), encyclopedic ($E_{C}$), logical ($L_{C}$), and tautological ($L_{T}$) sets.
@@ 4.1 Reasoning Format
The informal reasoning method instructs an LLM to answer $q$ with Yes or No based on a context. We employ few-shot prompting using three manually engineered in-context examples. To make the model more robust, two of the three in-context examples include distractions with irrelevant information. We explicitly note in the instruction part of the prompt that the provided contexts may contain irrelevant details. The model is evaluated in two modes: direct prompting, where it answers directly, and CoT prompting, where it generates step-by-step reasoning in natural language before generating an answer. To support CoT prompting, we manually created in-context examples with fine-grained natural language reasoning steps designed to improve performance Wei et al. (2022). The model’s answers are extracted using a regular expression (details in Appendix E).
We include an autoformalisation method combining a symbolic theorem prover with an LLM. This approach follows Logic-LM Pan et al. (2023) to formalise the context and query into symbolic representations, which are then deterministically evaluated. The process is divided into two subtasks: First, the model generates formal representations of the context and query, referred to as $c_{LF}$ (context logical form) and $q_{LF}$ (query logical form). Second, a symbolic theorem prover evaluates these logical forms to determine whether $q_{LF}$ can be derived from $c_{LF}$, producing a true or false outcome. This approach allows for a transparent and verifiable reasoning process grounded in logical consistency. In practice, we prompt the LLM to create the logical forms using the same three in-context examples as in the informal reasoning approach. The prompt is extended with instructions describing the formalisation syntax, following the methodology of Logic-LM. The resulting logical forms are parsed and combined into an Abstract Syntax Tree (AST), which provides the input for a theorem prover.
@@ 4.2 Formalisation Syntax
While the reasoning performance should be independent of the particular formalisation syntax, models may perform better with specific syntaxes, e.g., because of their frequency in the training data Razeghi et al. (2022). It is an open question whether the choice of syntax for formal reasoning impacts the model’s translation performance and the robustness of its reasoning. Although all the syntaxes we consider represent First-Order Logic (FOL), their surface form variations may influence the formalisation ability of LLMs. The three evaluated syntaxes, illustrated in Figure 3, contain identical information and are interchangeable, which ensures flexibility and allows the framework to include other syntaxes in the future:
@@
is widely used in logic classes and academic papers. It incorporates mathematical symbols such as $\forall$ and $\exists$ and implicitly distinguishes between variables and individuals. This syntax is employed by Logic-LM Pan et al. (2023).
@@
as the abbreviation of Thousands of Problems for Theorem Provers is a Prolog-like formalisation language developed for theorem provers. It avoids mathematical symbols and mandates that variables begin with an uppercase letter. While TPTP supports higher-order logic, we limit our scope to its first-order fragment (fof), using a syntax derived from its official specification Sutcliffe (2024).
If an individual drinks water, they will be hydrated. $\forall~{}x~{}drinkWater(x)\implies hydrated(x)$ FOL ?$x~{}drinkWater(?x)\implies hydrated(?x)$ R-FOL fof(a0,axiom,![X]:drinkWater(X) => hydrated(X)). TPTP
Figure 3: Examples of the three syntaxes: FOL, R-FOL, and TPTP.
@@ 4.3 Error Recovery Mechanism
The baseline approach does not attempt to correct errors. Instead, we predict a random value, true or false, as a fallback strategy. We avoid introducing an evaluation bias associated with more complex strategies, such as CoT-based refinement, which is commonly done in prior work Pan et al. (2023); Kirtania et al. (2024), as this would blur the comparison with other methods.
@@
The LLM is prompted to refine the logical form using a generic parsing error message, such as ’parsing error’. This type of message does not need a parser with an error handler, though it fails to point to specific errors in the logical form. Prior work has shown the effectiveness of this simple feedback method Pan et al. (2023).
@@
A more detailed approach where the LLM is given specific feedback, highlighting the exact parts of the logical form that violate the syntax. Creating this kind of feedback necessitates an error-handling strategy for the parser. For example, the missing argument in $man\wedge mortal(Socrates)$ results in the error message: mismatched input $\wedge$’ expecting ’(’. Using error messages from parsers as feedback to improve LLMs performance when synthesizing a formal language has shown promising results in code generation Zhang et al. (2023); Jiang et al. (2024).
@@
This strategy extends the error message with warnings generated from heuristics to recognize semantic errors, inspired by the “soft” critics in the LLM-modulo method Kambhampati et al. (2024), where it has been shown to enhance robustness. Notably, soft critics have not been incorporated in prior LLM-based methods for logical deduction.
@@ 5 Experimental Setup
The robustness evaluation is based on the FOL part of the LogicBench (Eval) Parmar et al. (2023) dataset, containing $520$ samples with $180$ unique contexts. LogicBench systematically covers nine different inference rules. The samples are automatically generated by prompting GPT-3.5 and manually verified. $340$ of the total of $520$ samples are negative examples constructed by negating the conclusion. If the conclusion is a disjunction, each part of the disjunction is negated, resulting in a slight class imbalance. The authors report a mean accuracy of around $90\%$ for three human annotators. By applying our robustness framework from §3, we obtain seven perturbed variants of LogicBench.
@@
The noise sentences are randomly sampled from a source $s$. We sample encyclopedic perturbation sentences from $10,000$ abstracts of Wikipedia articles gathered via its API. As a logical reasoning source, we use sentences from $1001$ contexts of the deduction QA benchmark FOLIO Han et al. (2022). As tautologies, we manually write $22$ sentences. All sentences use negations and, at most, one disjunction or conjunction. A complete list can be found in Appendix B. We randomly sample noise perturbations and add them to each sample. We do not alter the class distribution, i.e., we keep LogicBench’s original class imbalance.
We consider unique contexts for eight out of the nine logical forms to create counterfactual statement perturbations, resulting in $160$ samples. We create a balanced dataset by altering between valid and invalid queries from LogicBench.
@@
We test GPT 4o-mini, as well as a smaller and a larger variant for the three open-source LLM families: Gemma-2 (9b and 27b), Mistral (7b and Small), and Llama 3.1 (8b and 70b).
@@ 6 Evaluation
We provide three sets of insights into this section, organised as findings (F*). We quantitatively study the effect of the adversarial and counterfactual perturbations on the performance of informal reasoners and autoformalisation methods. Then, we dive deeper into method variants. Finally, we analyse the nature of formalisation errors made by the models.
@@ 6.1 Robustness Analysis
F1: Noise perturbations have a stronger effect on formalisation methods than informal LLM reasoners.
Table 1 shows that, on average, the accuracy of both direct and CoT informal reasoning remains between $73\%$ and $74\%$ in the face of added noise. While the autoformalisation method performs similarly to informal reasoners on the original dataset, its performance decreases between $4\%$ and $11\%$. The accuracy drops especially with logical (L) and tautological (T) distractions, whose logical language formats trick the LLM into formalizing the noisy clauses. On the other hand, the linguistically complex and more natural sentences of encyclopedic distractions show a minor effect, suggesting that LLMs successfully avoids formalizing the more complicated sentences.
@@
Table 1 shows that counterfactual statements cause a significant decrease in performance for both the informal reasoners and autoformalisation methods of between $12\%$ and $13\%$ on average. Moreover, this observation also holds for all tested models, i.e., none are robust towards counterfactual perturbations across every evaluated dimension. Even the strongest model, GPT 4o-mini, yields a performance of 63-68%, which is relatively close to the random performance of 50%. The high impact of counterfactual statements (the single “not” inserted) could be due to the inability of LLMs to overwrite prior knowledge with explicitly stated information or memorization of the answers. We study the error sources further in §6.3.
@@ 6.2 Impact of Method Design
The evaluation of different logical forms for formal LLM-based reasoning in Table 2 shows the preference of some models for specific syntactic formats. Llama 3.1 70B has a considerable improvement of $12\%$ with TPTP syntax on the original set, while Llama 3.1 8B benefits from the R-FOL syntax. However, all grammars show a declining accuracy trend and increased syntax errors for noise perturbations, where the best grammar loses its advantage over the rest. When comparing the grammars on the counterfactual partitions, we observe that TPTP is consistently more robust than the standard first-order logic grammar. Here, GPT 4o-mini shows a reduction from $O$ to $O_{C}$ of $20\%$ for FOL and only $12\%$ for the TPTP grammar. Since this does not correlate with fewer syntax errors, the formalisation in TPTP prevents semantical errors for counterfactual premises. A positive reading of these results, especially the minor differences between FOL and R-FOL, is that autoformalisation LLMs can adapt to the grammar syntax prescribed in the prompt without further loss in performance.
@@
Table 3 shows the results with different error recovery mechanisms. The results indicate that no feedback strategy emerges as a winner in the different datasets. All feedback variants reduce syntax errors for noise perturbations, but given the lack of a consistent increase in accuracy, the corrected formalisations are most likely to contain semantic errors still. The type of feedback message only has a minor influence on correcting syntax errors, whereas Llama 3.1 70b and GPT 4o-mini correct slightly more syntax errors with specific error messages. This finding aligns with Huang et al. (2023), who also found that LLMs cannot consistently self-correct their reasoning after receiving relevant feedback.
@@ 6.3 Error Analysis
The low performance for noise perturbations correlates with more syntax errors for all models and distraction categories (cf. execution rates in Table 4). The three worst-performing models (both Mistral models, Gemma-2 9b) generate, at best, for $37\%$ and, at worst, for only $4\%$ of the samples, a valid logical form. Gemma-2 9b and Llama3.1 8b produce more syntax errors than the larger counterparts, suggesting that larger models are more robust towards noise perturbations. The accuracy of syntactically valid samples is higher than the informal reasoning methods for most distractions (Table 5), motivating informal reasoning as a backup strategy for formal reasoning. The error message feedback reveals two common syntax errors: 1) errors by models with an initial low execution rate exhibit issues with the template structure, including using incorrect keywords or adding conversational phrases; 2) perturbation-related errors, the most common of which is using undefined truth constants as part of tautological distractions.
@@
Unlike the introduced noise, counterfactual perturbations do not lead to more syntax errors. The execution rate in Table 4 is stable or improves for counterfactuals. However, we see a drop in accuracy for the counterfactual column $\text{O}_{C}$ in Table 1 and can conclude that the number of logical forms with semantic errors has to increase. This suggests that the introduced negation is not correctly formalised. Looking at the warnings generated by the feedback mechanism, for GPT 4o-mini, $161$ warning messages are generated on the unperturbed data. $54$ of these were fixed with a single iteration. Not considering predicates and individuals as part of the context is the most frequent warning across all models.

### v2

@@ 3.2 Counterfactual Shifts
Formalisation. We consider common-sense contradictions by altering original statements in $c$ into counterfactual ones. The inclusion of counterfactual statements is motivated by the requirement for faithful reasoning. Namely, the validity of a deduction depends only on the structure of the logical form, which in turn should follow the original task description in the context and the question. Therefore, the reasoning capabilities of a model should not rely on prior knowledge. Instead, explicitly stated premises should be prioritised over implicit model knowledge and overwrite it. This behaviour allows models to reason over hypothetical and counterfactual premises, enabling the derivation of new (scientific) knowledge, e.g., deriving non-Euclidean geometry by replacing the parallel postulate, or science fiction authors to create complete, logically consistent alternative worlds based on one changed premise. Moreover, it ensures a more controlled evaluation setup for evaluating reasoning independently of the quality of the internal model knowledge, since all models use the same premises. To introduce counterfactual premises, we do not add new sentences; instead, we negate sentences from the original context $c$. Since original premises often state common-sense knowledge, their negation naturally contradicts world knowledge. For example, in the modus ponens inference in Figure 1, we negate $mortal(X)$, stating that All men are immortal. The altered context $c^{\prime}$ is formalised as follows: $c^{\prime}=neg(c)$, where $neg(c)$ negates one of the terms in the original context. Naturally, these perturbations influence the inferred answer and depend on the applied inference rule. The negated term is selected so that the label of the resulting function $f^{\prime}$ is opposite from that of the original function $f$.
Design of counterfactual perturbations. We assume a deductive reasoning dataset where the natural language form corresponds to a logical formula. Then, we introduce counterfactual perturbations by altering each logical formula using predefined rules that negate the formula (see the full list of rules in the Appendix). For example, we negate the consequent $\mathsf{q}(a)$ for the first implication of a constructive dilemma and adapt the inference to $\neg\mathsf{q}(a)\vee\mathsf{s}(a)$. Then, the negation to create the context $c^{\prime}$ is manually added before the relevant terms in $c$, thus guaranteeing high data quality. Since the inference results in the opposite label for $f^{\prime}$, we adapt the target label accordingly. Figure 2 (top-right) shows an example of such a contradiction. Importantly, all counterfactual perturbations can be combined with noise perturbations, leading to counterfactual versions of the original ($O_{C}$), encyclopedic ($E_{C}$), logical ($L_{C}$), and tautological ($L_{T}$) sets.
@@ 4.1 Reasoning Format
The informal reasoning method instructs an LLM to answer $q$ with Yes or No based on a context. We employ few-shot prompting using three manually engineered in-context examples. To make the model more robust, two of the three in-context examples include distractions with irrelevant information. We explicitly note in the instruction part of the prompt that the provided contexts may contain irrelevant details. The model is evaluated in two modes: direct prompting, where it answers directly, and CoT prompting, where it generates step-by-step reasoning in natural language before generating an answer. To support CoT prompting, we manually created in-context examples with fine-grained natural language reasoning steps designed to improve performance [32]. The model’s answers are extracted using a regular expression (details in Appendix D).
We include an autoformalisation method combining a symbolic theorem prover with an LLM. This approach follows Logic-LM [18] to formalise the context and query into symbolic representations, which are then automatically proven. The process is divided into two subtasks: First, the model generates formal representations of the context and query, referred to as $c_{LF}$ (context logical form) and $q_{LF}$ (query logical form). Second, a symbolic theorem prover evaluates these logical forms to determine whether $q_{LF}$ can be derived from $c_{LF}$, producing a true or false outcome. This approach allows for a transparent and verifiable reasoning process grounded in logical consistency. In practice, we prompt the LLM to create the logical forms using the same three in-context examples as in the informal reasoning approach. The prompt is extended with instructions describing the formalisation syntax, following the methodology of Logic-LM. The resulting logical forms are parsed and combined into an Abstract Syntax Tree (AST), which provides the input for a theorem prover.
@@ 4.2 Formalisation Syntax
While the reasoning performance should be independent of the particular formalisation syntax, models may perform better with specific syntaxes, e.g., because of their frequency in the training data [21]. It is an open question whether the choice of syntax for formal reasoning impacts the model’s translation performance and the robustness of its reasoning. Although all the syntaxes we consider represent First-Order Logic (FOL), their surface form variations may influence the formalisation ability of LLMs. The three evaluated syntaxes, illustrated in Figure 3, contain identical information and are interchangeable, which ensures flexibility and allows the framework to include other syntaxes in the future:
@@
is widely used in logic classes and academic papers. It incorporates mathematical symbols such as $\forall$ and $\exists$ and implicitly distinguishes between variables and individuals. This syntax is employed by Logic-LM [18].
@@
as the abbreviation of Thousands of Problems for Theorem Provers is a Prolog-like formalisation language developed for theorem provers. It avoids mathematical symbols and mandates that variables begin with an uppercase letter. While TPTP supports higher-order logic, we limit our scope to its first-order fragment (fof), using a syntax derived from its official specification [26].
If an individual drinks water, they will be hydrated. $\forall~x~drinkWater(x)\implies hydrated(x)$ FOL $\forall~?x~drinkWater(?x)\implies hydrated(?x)$ R-FOL fof(a0,axiom,![X]:drinkWater(X) => hydrated(X)). TPTP
Figure 3: Examples of the three syntaxes: FOL, R-FOL, and $TP$.
@@ 4.3 Error Recovery Mechanism
The baseline approach does not attempt to correct errors. Instead, we predict a random value, true or false, as a fallback strategy. We avoid introducing an evaluation bias associated with more complex strategies, such as CoT-based refinement, which is commonly done in prior work [18, 12], as this would blur the comparison with other methods.
@@
The LLM is prompted to refine the logical form using a generic parsing error message, such as ’parsing error’. This type of message does not need a parser with an error handler, though it fails to point to specific errors in the logical form. Prior work has shown the effectiveness of this feedback method [18].
@@
A more detailed approach where the LLM is given specific feedback, highlighting the exact parts of the logical form that violate the syntax. Creating this kind of feedback necessitates an error-handling strategy for the parser. For example, the missing argument in $man\wedge mortal(Socrates)$ results in the error message: mismatched input $\wedge$’ expecting ’(’. Using error messages from parsers as feedback to improve LLMs performance has shown promising results in code generation [35, 10].
@@
An extension of the error messages with warnings generated from heuristics to recognise semantic errors, inspired by the “soft” critics in the LLM-modulo method [11], where it has been shown to enhance robustness. Notably, soft critics have not been incorporated in prior LLM-based methods for logical deduction.
@@ 5 Experimental Setup
The experimental setup, including the generated perturbed data, the source code and configuration files, is made available on Zenodo [8] and GitHub [^1] https://github.com/Fab-Hop/langdeductive to facilitate reproducibility.
@@
The robustness evaluation is based on the FOL part of the LogicBench (Eval) [19] dataset, containing $520$ samples with $180$ unique contexts. LogicBench systematically covers nine different inference rules. The samples are automatically generated by prompting GPT-3.5 and manually verified. $340$ of the total of $520$ samples are negative examples constructed by negating the conclusion. If the conclusion is a disjunction, each part of the disjunction is negated, resulting in a slight class imbalance. The authors report a mean accuracy of around $90\%$ for three human annotators. By applying our robustness framework from §3, we obtain seven perturbed variants of LogicBench.
@@
The noise sentences are randomly sampled from a source $s$. We sample encyclopedic perturbation sentences from $10,000$ abstracts of Wikipedia articles gathered via its API. As a logical reasoning source, we use sentences from $1001$ contexts of the deduction QA benchmark FOLIO [6]. As tautologies, we manually write $22$ sentences. All sentences use negations and, at most, one disjunction or conjunction. A complete list can be found in Appendix B. We randomly sample noise perturbations and add them to each sample. We do not alter the class distribution, i.e., we keep LogicBench’s original class imbalance. The three types of noise sources have varying, yet consistently low, levels of semantic relevance, quantified as an average cosine similarity of sentence-transformers embeddings between the added noise and existing context. Tautological noise is most similar ($0.20$), followed by logical noise ($0.09$), and encyclopedic noise ($0.04$). The low semantic similarity and the careful construction of the dataset, i.e. appending distractions only at the beginning of the sentence ensure no interference with the original semantics.
We consider unique contexts for eight out of the nine logical forms to create counterfactual statement perturbations, resulting in $160$ samples. We create a balanced dataset by alternating between valid and invalid queries from LogicBench. This dataset is a synthetic dataset created by an LLM using its latent knowledge. Thus, the original premises align with world knowledge encoded in an LLM and the negation of a premise creates a counterfactual premise adapted to this latent knowledge.
@@
We test GPT 4o-mini, as well as a smaller and a larger variant for the three open-source LLM families: Gemma-2 (9b and 27b), Mistral (7b and Small), and Llama 3.1 (8b and 70b). If available, the tested LLMs are used in their instruction-tuned variant. Except for GPT 4o-mini, all models are provided from hugging-face and used with 4-bit quantisation. Our experiments use a default temperature of $1.0$ to keep the output distributions unmodified. Furthermore, we verify the consistency of our findings by conducting experiments with a temperature setting of 0.
Syntax Parsers.
For each grammar in § 4.2, we generate a parser using ANTLR4222https://www.antlr.org/ to decouple the logical forms from the input requirements of the theorem prover. The FOL variant is adapted from the ANTLR grammars-v4 repository [^3] https://github.com/antlr/grammars-v4. The parser validates the syntax and, if necessary, translates it into TPTP to pass on to the theorem prover. The Vampire theorem prover [^4] https://vprover.github.io processes this formal method’s input, generating proofs by refutation to check whether the query logically follows from the context.
The syntax parser generates specific feedback for all error recovery strategies to prompt the LLM. The LLM prompt includes three in-context examples. When recovery is attempted, refinement is limited to three iterations. If the error persists, the random fallback strategy is applied. This setup allows for systematic comparison of the robustness of these recovery methods in the autoformalisation process. The generated warnings use three heuristics to identify semantic errors. First, checking for predicates and individuals only mentioned in the query to avoid an incomplete context. Second, ensuring that all predicates with the same identifier use the same number of parameters (checking for n-arity). Finally, avoiding similarly named predicates and individuals based on Levenshtein distance.
@@ 6 Evaluation
We provide three sets of insights into this section, organised as findings (F*). We quantitatively study the effect of adversarial and counterfactual perturbations on the performance of informal reasoners and autoformalisation methods. Then, we dive deeper into method variants. Finally, we analyse the nature of formalisation errors made by the models.
@@ 6.1 Robustness Analysis
F1: Noise perturbations have a stronger effect on autoformalisation methods than informal LLM reasoners.
Table 1 shows that, on average, the accuracy of both direct and CoT informal reasoning remains between $73\%$ and $74\%$ in the face of added noise. While the autoformalisation method performs similarly to informal reasoners on the original dataset, its performance decreases between $4\%$ and $11\%$. The accuracy drops especially with logical (L) and tautological (T) distractions, whose language formats trick the LLM into formalising the noisy clauses. On the other hand, the linguistically complex and more natural sentences of encyclopedic distractions show a minor effect, suggesting that LLMs successfully avoids formalising the more complicated but semantically less similar noise sentences.
@@
Table 1 shows that counterfactual statements cause a significant decrease in performance for both the informal reasoners and autoformalisation methods of between $12\%$ and $13\%$ on average. Moreover, this observation also holds for all tested models, i.e., none are robust towards counterfactual perturbations across every evaluated dimension. Even the strongest model, GPT 4o-mini, yields a performance of 63-68%, which is relatively close to the random performance of 50%. The high impact of counterfactual statements (the single “not” inserted) could be due to the inability of LLMs to overwrite prior knowledge with explicitly stated information or memorisation of the answers. We study the error sources further in §6.3.
@@ 6.2 Impact of Method Design
The evaluation of different logical forms for formal LLM-based reasoning in Table 2 shows the preference of some models for specific syntactic formats. Llama 3.1 70B has a considerable improvement of $12\%$ with TPTP syntax on the original set, while Llama 3.1 8B benefits from the R-FOL syntax. However, all grammars show a declining accuracy trend and increased syntax errors for noise perturbations, where the best grammar loses its advantage over the rest. When comparing the grammars on the counterfactual partitions, we observe that TPTP is consistently more robust than the standard first-order logic grammar. Here, GPT 4o-mini shows a reduction from $O$ to $O_{C}$ of $20\%$ for FOL and only $12\%$ for the TPTP grammar. Since this does not correlate with fewer syntax errors, the formalisation in TPTP prevents semantic errors for counterfactual premises. A positive reading of these results, especially the minor differences between FOL and R-FOL, is that autoformalisation LLMs can adapt to the grammar syntax prescribed in the prompt without further loss in performance.
@@
Table 3 shows the results with different error recovery mechanisms. The results indicate that no feedback strategy emerges as a winner in the various datasets. All feedback variants reduce syntax errors for noise perturbations, but given the lack of a consistent increase in accuracy, the corrected formalisations are still most likely to contain semantic errors. The type of feedback message only has a minor influence on correcting syntax errors, whereas Llama 3.1 70b and GPT 4o-mini correct slightly more syntax errors with specific error messages. This finding aligns with [9], who also found that LLMs cannot consistently self-correct their reasoning after receiving relevant feedback.
@@ 6.3 Error Analysis
The low performance for noise perturbations correlates with more syntax errors for all models and distraction categories (cf. execution rates in Table 4). The three worst-performing models (both Mistral models, Gemma-2 9b) generate, at best, for $37\%$ and, at worst, for only $4\%$ of the samples, a valid logical form. In particular, these models show an initial low execution rate between $41\%$ and $53\%$. The error message feedback reveals issues with the template structure, including using incorrect keywords or adding conversational phrases. Noise perturbations emphasize this effect causing the increased syntax errors. The second common syntax errors are perturbation-related. The added noise triggers the LLMs to hallucinate new syntax rules, e.g., introducing undefined truth constants as part of tautological distractions or switching from a camelCase to snake_case naming schema as part of logical distractions, which is likely caused by the increased difference between in-context examples and the test sample. The error analysis suggests that larger models are more robust towards noise perturbations. Gemma-2 9b and Llama3.1 8b produce more syntax errors than their larger counterparts.
As the increased syntax errors are mainly consistent for specific noise types, e.g., tautological noise creates undefined truth constants, it may be suboptimal to use a single, larger LLM for a broad set of formalisation tasks. Instead a framework with specialized agents that address particular syntax errors could mitigate these errors more effectively. Additionally, the accuracy of syntactically valid samples is higher than the informal reasoning methods for most distractions (Table 5), motivating a hybrid strategy: use formal reasoning when a valid formalisation is produced, and otherwise rely on informal reasoning as a backup.
@@
Unlike the introduced noise, counterfactual perturbations do not lead to more syntax errors. The execution rate in Table 4 is stable or improves for counterfactuals. However, we see a drop in accuracy for the counterfactual column $\text{O}_{C}$ in Table 1 and can conclude that the number of logical forms with semantic errors has to increase. This suggests that the introduced negation is not correctly formalised with respect to the semantics of these samples. Instead of genuine deductive reasoning capabilities informal and autoformalisation rely on prior knowledge.
Devising principled mechanisms for verifying semantic validity (e.g., ensuring that predicates are defined before being used in a query) could reduce semantic errors for counterfactuals. Looking at the warnings generated by the feedback mechanism, for GPT 4o-mini, $161$ warning messages are generated on the unperturbed data pointing towards specific semantic errors. $54$ of these were fixed with a single iteration. Across all models, the most frequent warning is not considering predicates and individuals as part of the context. This suggest that verification feedback is able to verify semantic validity and incorporating this feedback can mitigate semantic errors.